# 1st Runze Dong    2nd Buhong Wang    3rd Cunqian Feng    4th Jiang Weng    5th Chen Han    6th Jiwei Tian

### v1

@@ I Introduction
With the growing demand for ubiquitous connectivity, immersive communication, and high spectrum utilization and potential application scenarios such as smart city, industrial automation, and autonomous driving for the next generation communication networks, integrated sensing and communication (ISAC) has brought about widespread attention [1, 2]. This integration is particularly critical in dynamic and complex scenarios where efficient resource allocation and real-time adaptability are significant. Unmanned aerial vehicles (UAVs) are renowned for their inherent mobility, rapid deployment capability, and favorable line-of-sight (LoS) propagation conditions. These qualities have led to the emergence of UAVs as flexible and agile platforms capable of enhancing such ISAC-enabled networks, which makes them ideal candidates for applications ranging from coverage expansion and public safety to precision monitoring and disaster response. Therefore, the convergence of ISAC and UAV represents a promising path towards next-generation wireless systems that are both spectral efficient and event aware [3, 4, 5].
ISAC enabled UAV systems have garnered significant research attention, with prior work focusing on trajectory and beamforming co-design [6] and resource allocation under performance trade-offs [5]. Despite these advances, few works focus on the security performance [7, 8]. An intelligent reflecting surface (IRS) was introduced in [7] to assist the secure transmission of UAV, and its transmit power, user scheduling, and trajectory were jointly optimized to maximize the average achievable rate. In [8], alternating optimization (AO) strategy was employed to design the trajectory and beamforming of an UAV base station.
@@
The increasing complexity and capacity demands of wireless networks have positioned DRL as a critical tool for addressing non-convex optimization challenges in dynamic communication networks. An millimeter wave (mmWave) heterogeneous network including diverse multiple access schemes was considered in [9], and DRL was employed to tackle a multi-objective optimization problem aimed at maximizing user fairness while minimizing transmit power. Similarly, [10] investigated an mmWave vehicular network scenario where DRL was leveraged to enhance the secrecy capacity of all target vehicles under energy consumption constraint. Furthermore, with the emergence of artificial intelligence generated content (AIGC) services, multiple UAVs were integrated as agile service providers in [11], and a diffusion model-enhanced DRL framework was developed to optimize computation task offloading and resource management policies. These studies collectively demonstrate the adaptability and effectiveness of DRL in handling complex, high-dimensional optimization problems across various wireless network architectures.
Motivated by the aforementioned challenges and opportunities, this paper proposes a joint beamforming and AN design to simultaneously enhance communication security and spectral efficiency in dynamic multi-UAV ISAC networks. As the pricing budget for configuring each antenna with a dedicated radio frequency (RF) chain, the hybrid beamforming (HBF) framework is adopted [12]. The objective is to maximize the sum secrecy rate across all legitimate UAVs, by jointly optimizing the hybrid beamformer, AN, and trajectories of legitimate UAVs. A two-stage optimization framework is developed to balance performance and computational tractability. In the first stage, the digital beamformer and UAV trajectories are optimized using Proximal Policy Optimization (PPO), chosen for its favorable convergence properties and training stability. The second stage employs an AO-based method with low-complexity to decompose the fully-digital solution into its analog and digital components under power and constant modulus constraints.
@@ II-B Communication Model
where $\bm{f}_{BB,l}[n]\in\mathbb{C}^{N_{RF}\times 1}$ is the digital beamformer for signal $s_{l}[n]$. To facilitate comprehension, define $\bm{F}_{BB}[n]\triangleq\left[\bm{s}_{1}[n],\bm{s}_{2}[n],\ldots,\bm{s}_{L}[n]\right]$ as the overall digital beamformer for all confidential signals, which will leading (1) to
@@
in which $\bm{s}_{n}=\left[s_{1}[n],s_{2}[n],\ldots,s_{L}[n]\right]^{T}$ denoting the $L$ confidential data streams corresponding to $L$ legitimate UAVs, $z[n]$ is the AN and $\bm{w}[n]\in\mathbb{C}^{N_{RF}\times 1}$ is the digital beamformer for it. It is assumed that the confidential signal $\bm{s}[n]$ and AN are statistically independent, with $\mathbb{E}[\bm{s}[n]\bm{s}^{H}[n]]=\bm{I}_{L}$ and $\mathbb{E}[|z[n]|^{2}]=1$. $\bm{F}_{RF}[n]\in\mathbb{C}^{N_{t}\times N_{RF}}$ is the analog beamformer. As such, the received signals at the $l$-th legitimate UAV and the $e$-th eavesdropping UAV could be expressed as
$y_{l}[n]=\bm{h}_{l}^{H}[n]\bm{x}[n]+n_{l}[n],\vskip-3.00003pt$ (3a)
$y_{e}[n]=\bm{h}_{e}^{H}[n]\bm{x}[n]+n_{e}[n],\vskip-3.00003pt$ (3b)
where $\bm{h}_{l}[n]\in\mathbb{C}^{N_{t}\times 1}$ and $\bm{h}_{e}[n]\in\mathbb{C}^{N_{t}\times 1}$ represent the channel gains from the ISAC base station to the $l$-th legitimate UAV and the $e$-th eavesdropper UAV respectively. The additive white Gaussian noises (AWGN) at the corresponding receivers is denoted by $n_{l}[n]\sim\mathcal{CN}(0,\sigma_{l}^{2})$, $n_{e}[n]\sim\mathcal{CN}(0,\sigma_{e}^{2})$. Following the simplified mmWave channel model [13], the channel gains $\bm{h}_{i},i\in\{l,e\}$ is given by
$\bm{h}_{i}=\beta_{0}\alpha_{0}\mathbf{a}(\phi_{i},\theta_{i}),\vskip-3.00003pt$ (4)
where $\lambda_{0}=\frac{c}{4\pi f_{c}}d_{i}^{-\kappa}$ is the channel gain coefficients, $d_{i}$ is the distance between the ISAC base station and node $i$, $\alpha_{0}\sim\mathcal{CN}(0,1)$ is the small-scale fading gain, $\mathbf{a}(\phi,\theta)$ is the steering vector corresponding to the UPA on the ISAC base station
@@
$\begin{split}y_{l}[n]\!&=\!\underbrace{\bm{h}_{l}^{H}\bm{F}_{RF}[n]\bm{f}_{BB,l}[n]\bm{s}_{l}[n]}_{\text{Desired Signal}}\!+\!\underbrace{\sum_{j\neq l}^{L}\bm{h}_{l}^{H}\bm{F}_{RF}[n]\bm{f}_{BB,j}[n]\bm{s}_{j}[n]}_{\text{Multi-User Interference}}\\ &+\underbrace{\bm{h}_{l}^{H}\bm{F}_{RF}[n]\bm{w}[n]\bm{z}[n]}_{\text{AN Interference}}+n_{l}[n],\end{split}$ (6a)
$\begin{split}y_{e,l}[n]\!&=\!\underbrace{\bm{h}_{e}^{H}\bm{F}_{RF}[n]\bm{f}_{BB,l}[n]\bm{s}_{l}[n]}_{\text{Signal in Interest}}\!+\!\underbrace{\sum_{j\neq l}^{L}\bm{h}_{e}^{H}\bm{F}_{RF}[n]\bm{f}_{BB,j}[n]\bm{s}_{j}[n]}_{\text{Interference}}\\ &+\underbrace{\bm{h}_{e}^{H}\bm{F}_{RF}[n]\bm{w}[n]\bm{z}[n]}_{\text{Jamming by AN}}+n_{e}[n],\vskip-3.00003pt\end{split}$ (6b)
and the received signal-to-interference ratio (SINR) could be given by
@@ IV-A3 Reward $\mathcal{R}$
The total reward consists three parts as $R_{n}=R_{com}[n]+R_{sen}[n]+R_{QoS}[n]$. $R_{com}[n]$ is the secure communication part, which equals the sum secrecy rate of all legitimate UAVs at step $n$. $R_{sen}$ and $R_{QoS}$ are the sensing performance penalty and QoS penalty respectively
$R_{sen}=-\textstyle\sum_{e=1}^{E}\left[\Gamma_{e}-P_{\phi_{,}\theta_{e}}d_{e}^{-2}\right]^{+},\vskip-3.00003pt$ (15a)
$R_{QoS}=-\textstyle\sum_{l=1}^{L}\left[R_{min}-R_{s,l}[n]\right]^{+}.\vskip-3.00003pt$ (15b)
@@ V Simulation Experiments
The ISAC base station is assumed to be located at the origin, and the height of its antenna array is neglected. The considered multi-UAV ISAC network comprises $L=\text{4}$ legitimate UAVs, with initial locations at $\text{20},\text{80},\text{20}]^{T}$, $\text{20},\text{70},\text{15}]^{T}$, $\text{70},\text{30},\text{30}]^{T}$, and $\text{80},\text{10},\text{15}]^{T}$ respectively. Meanwhile, $E=\text{3}$ eavesdropper UAVs are positioned near the legitimate UAVs at locations $\text{20},\text{40},\text{20}]^{T}$, $\text{50},\text{50},\text{30}]^{T}$, and $\text{80},\text{70},\text{40}]^{T}$. The UPA consists $N_{t}=8\times 8$ antennas, and the number of RF chains is $N_{RF}=8$. The transmit power of ISAC base station is $P_{t}=10$\mathrm{W}$\$ and the loss component of channel gain is $\kappa=1.8$. With time interval $\Delta_{t}=0.5$\mathrm{s}$\$ and total transmission duration $T=100$\mathrm{s}$\$, we have the number of time slots $N=200$. Sensing performance threshold is set as $\Gamma_{e}=0.5e^{-5}$ and the sensing penalty is $\zeta=0.5$. For all UAVs the noise power is $\sigma_{i}^{2}=1e^{-13}$.

### v2

@@ I Introduction
With the growing demand for ubiquitous connectivity, immersive communication, and high spectrum utilization and potential application scenarios such as smart city, industrial automation, and autonomous driving for the next generation communication networks, integrated sensing and communication (ISAC) has brought about widespread attention [RN1139, RN1095]. This integration is particularly critical in dynamic and complex scenarios where efficient resource allocation and real-time adaptability are significant. Unmanned aerial vehicles (UAVs) are renowned for their inherent mobility, rapid deployment capability, and favorable line-of-sight (LoS) propagation conditions. These qualities have led to the emergence of UAVs as flexible and agile platforms capable of enhancing such ISAC-enabled networks, which makes them ideal candidates for applications ranging from coverage expansion and public safety to precision monitoring and disaster response. Therefore, the convergence of ISAC and UAV represents a promising path towards next-generation wireless systems that are both spectral efficient and event aware [RN1094, RN1087, RN1028].
ISAC enabled UAV systems have garnered significant research attention, with prior work focusing on trajectory and beamforming co-design [RN873] and resource allocation under performance trade-offs [RN1028]. Despite these advances, few works focus on the security performance [RN1029, RN1132]. An intelligent reflecting surface (IRS) was introduced in [RN1029] to assist the secure transmission of UAV, and its transmit power, user scheduling, and trajectory were jointly optimized to maximize the average achievable rate. In [RN1132], alternating optimization (AO) strategy was employed to design the trajectory and beamforming of an UAV base station.
@@
The increasing complexity and capacity demands of wireless networks have positioned DRL as a critical tool for addressing non-convex optimization challenges in dynamic communication networks. An millimeter wave (mmWave) heterogeneous network including diverse multiple access schemes was considered in [RN1144], and DRL was employed to tackle a multi-objective optimization problem aimed at maximizing user fairness while minimizing transmit power. Similarly, [RN1036] investigated an mmWave vehicular network scenario where DRL was leveraged to enhance the secrecy capacity of all target vehicles under energy consumption constraint. Furthermore, with the emergence of artificial intelligence generated content (AIGC) services, multiple UAVs were integrated as agile service providers in [RN1104], and a diffusion model-enhanced DRL framework was developed to optimize computation task offloading and resource management policies. These studies collectively demonstrate the adaptability and effectiveness of DRL in handling complex, high-dimensional optimization problems across various wireless network architectures.
Motivated by the aforementioned challenges and opportunities, this paper proposes a joint beamforming and AN design to simultaneously enhance communication security and spectral efficiency in dynamic multi-UAV ISAC networks. As the pricing budget for configuring each antenna with a dedicated radio frequency (RF) chain, the hybrid beamforming (HBF) framework is adopted [RN1142]. The objective is to maximize the sum secrecy rate across all legitimate UAVs, by jointly optimizing the hybrid beamformer, AN, and trajectories of legitimate UAVs. A two-stage optimization framework is developed to balance performance and computational tractability. In the first stage, the digital beamformer and UAV trajectories are optimized using Proximal Policy Optimization (PPO), chosen for its favorable convergence properties and training stability. The second stage employs an AO-based method with low-complexity to decompose the fully-digital solution into its analog and digital components under power and constant modulus constraints.
@@ II-B Communication Model
where $\bm{F}_{RF}[n]\in\mathbb{C}^{N_{t}\times N_{RF}}$ is the analog beamformer, $\bm{f}_{BB,l}[n]\in\mathbb{C}^{N_{RF}\times 1}$ is the digital beamformer for signal $s_{l}[n]$. To facilitate comprehension, define $\bm{F}_{BB}[n]\triangleq\left[\bm{f}_{BB,1}[n],\bm{f}_{BB,2}[n],\ldots,\bm{f}_{BB,L}[n]\right]$ as the overall digital beamformer for all confidential signals, which will leading (1) to
@@
in which $\bm{s}[n]=\left[s_{1}[n],s_{2}[n],\ldots,s_{L}[n]\right]^{T}$ denoting the $L$ confidential data streams corresponding to $L$ legitimate UAVs, $z[n]$ is the AN and $\bm{w}[n]\in\mathbb{C}^{N_{RF}\times 1}$ is the digital beamformer for it. It is assumed that the confidential signal $\bm{s}[n]$ and AN are statistically independent, with $\mathbb{E}[\bm{s}[n]\bm{s}^{H}[n]]=\bm{I}_{L}$ and $\mathbb{E}[|z[n]|^{2}]=1$. As such, the received signals at the $l$-th legitimate UAV and the $e$-th eavesdropping UAV could be expressed as
$y_{l}[n]=\bm{h}_{l}^{H}[n]\bm{x}[n]+n_{l}[n],$ (3a)
$y_{e}[n]=\bm{h}_{e}^{H}[n]\bm{x}[n]+n_{e}[n],$ (3b)
where $\bm{h}_{l}[n]\in\mathbb{C}^{N_{t}\times 1}$ and $\bm{h}_{e}[n]\in\mathbb{C}^{N_{t}\times 1}$ represent the channel gains from the ISAC base station to the $l$-th legitimate UAV and the $e$-th eavesdropper UAV respectively. The additive white Gaussian noises (AWGN) at the corresponding receivers is denoted by $n_{l}[n]\sim\mathcal{CN}(0,\sigma_{l}^{2})$, $n_{e}[n]\sim\mathcal{CN}(0,\sigma_{e}^{2})$. Following the simplified mmWave channel model [RN948], the channel gains $\bm{h}_{i},i\in\{l,e\}$ is given by
$\bm{h}_{i}=\beta_{0}\alpha_{0}\mathbf{a}(\phi_{i},\theta_{i}),\vskip-1.99997pt$ (4)
where $\beta_{0}=\frac{c}{4\pi f_{c}}d_{i}^{-\kappa}$ is the channel gain coefficients, $d_{i}$ is the distance between the ISAC base station and node $i$, $\alpha_{0}\sim\mathcal{CN}(0,1)$ is the small-scale fading gain, $\mathbf{a}(\phi,\theta)$ is the steering vector corresponding to the UPA on the ISAC base station
@@
$\begin{split}y_{l}[n]\!&=\!\underbrace{\bm{h}_{l}^{H}\bm{F}_{RF}[n]\bm{f}_{BB,l}[n]s_{l}[n]}_{\text{Desired Signal}}\!+\!\underbrace{\sum_{j\neq l}^{L}\bm{h}_{l}^{H}\bm{F}_{RF}[n]\bm{f}_{BB,j}[n]s_{j}[n]}_{\text{Multi-User Interference}}\\ &+\underbrace{\bm{h}_{l}^{H}\bm{F}_{RF}[n]\bm{w}[n]\bm{z}[n]}_{\text{AN Interference}}+n_{l}[n],\end{split}$ (6a)
$\begin{split}y_{e,l}[n]\!&=\!\underbrace{\bm{h}_{e}^{H}\bm{F}_{RF}[n]\bm{f}_{BB,l}[n]s_{l}[n]}_{\text{Signal in Interest}}\!+\!\underbrace{\sum_{j\neq l}^{L}\bm{h}_{e}^{H}\bm{F}_{RF}[n]\bm{f}_{BB,j}[n]s_{j}[n]}_{\text{Interference}}\\ &+\underbrace{\bm{h}_{e}^{H}\bm{F}_{RF}[n]\bm{w}[n]\bm{z}[n]}_{\text{Jamming by AN}}+n_{e}[n],\vskip-3.00003pt\end{split}$ (6b)
and the received signal-to-interference plus noise ratio (SINR) could be given by
@@ IV-A3 Reward $\mathcal{R}$
The total reward consists three parts as $R_{n}=R_{com}[n]+R_{sen}[n]+R_{QoS}[n]$. $R_{com}[n]$ is the secure communication part, which equals the sum secrecy rate of all legitimate UAVs at step $n$. $R_{sen}[n]$ and $R_{QoS}[n]$ are the sensing performance penalty and QoS penalty respectively
$R_{sen}[n]=-\textstyle\sum_{e=1}^{E}\left[\Gamma_{e}-P_{\phi_{,}\theta_{e}}[n]d_{e}^{-2}\right]^{+},\vskip-3.00003pt$ (15a)
$R_{QoS}[n]=-\textstyle\sum_{l=1}^{L}\left[R_{min}-R_{s,l}[n]\right]^{+}.\vskip-3.00003pt$ (15b)
@@ V Simulation Experiments
The ISAC base station is assumed to be located at the origin, and the height of its antenna array is neglected. The considered multi-UAV ISAC network comprises $L=\text{4}$ legitimate UAVs, with initial locations at $\text{20},\text{80},\text{20}]^{T}$, $\text{20},\text{70},\text{15}]^{T}$, $\text{70},\text{30},\text{30}]^{T}$, and $\text{80},\text{10},\text{15}]^{T}$ respectively. Meanwhile, $E=\text{3}$ eavesdropper UAVs are positioned near the legitimate UAVs at locations $\text{20},\text{40},\text{20}]^{T}$, $\text{50},\text{50},\text{30}]^{T}$, and $\text{80},\text{70},\text{40}]^{T}$. The UPA consists $N_{t}=8\times 8$ antennas, and the number of RF chains is $N_{RF}=8$. The transmit power of ISAC base station is $P_{t}=10$\mathrm{W}$\$ and the loss component of channel gain is $\kappa=1.8$. With time interval $\Delta_{t}=0.5$\mathrm{s}$\$ and total transmission duration $T=100$\mathrm{s}$\$, we have the number of time slots $N=200$. Sensing performance threshold is set as $\Gamma_{e}=0.5e^{-5}$ and the QoS requirement $R_{min}$ is 0.5 bps/Hz. For all UAVs the noise power is $\sigma_{i}^{2}=1e^{-13}$.